# The Race between Technology and Woman:

### v8

@@ 4.3 Factor substitution
Manufacturing 0.765 1.048 0.880 0.965 0.582 0.996 1.309 1.306 0.548 0.891 1.180 1.183
(0.152) (0.164) (0.185) (0.160) {0.413, 0.836} {0.894, $1$121, 1.589} {1.151, 1.540} {0.401, $0$753, 1.050} {0.955, 1.461} {0.990, 1.415}
Utilities 0.281 0.735 0.656 0.968 0.455 0.947 1.120 1.266 0.457 0.877 1.037 1.157
(0.278) (0.113) (0.160) (0.161) {0.306, 0.649} {0.697, 1.260} {0.898, 1.429} {1.038, 1.589} {0.319, 0.629} {0.603, 1.242} {0.775, 1.386} {0.891, 1.503}
Construction 0.637 0.832 0.775 0.663 1.057 1.402 1.453 1.517 0.888 1.176 1.289 1.342
(0.328) (0.447) (0.143) (0.169) {0.339, 2.544} {0.720, 2.602} {0.856, 2.375} {0.946, 2.415} {0.488, 1.508} {0.749, 1.789} {0.848, 1.932} {0.920, 1.971}
Services 0.691 0.905 0.788 0.920 0.737 0.989 1.214 1.310 0.689 0.890 1.104 1.167
(0.189) (0.100) (0.094) (0.079) {0.583, 0.919} {0.860, 1.184} {1.060, 1.429} {1.147, 1.569} {0.564, 0.814} {0.772, 1.019} {0.978, 1.249} {1.035, 1.322}
Notes: Estimates of $\epsilon_{f,n}$ in equation (27) for female skilled labor ($f=fh$), male skilled labor ($f=mh$), female unskilled labor ($f=fu$), and male unskilled labor ($f=mu$) are reported. Standard errors are in parentheses of the first four columns. The 95% Anderson–Rubin confidence sets are in curly brackets of the last eight columns. Both of them are robust to clustering at the country level. Lagged population size and lagged ICT investment price are used as instrumental variables in the fifth to eighth columns and the ninth to the last columns, respectively.
@@ 5 Estimation
This section describes the specification, identification, estimation, and aggregation of the sectoral production function.
@@ 5.1 Specification
The economy consists of multiple sectors. In each sector, production requires two types of capital (ICT capital, $k_{i}$, and non-ICT capital, $k_{o}$) and four types of labor (male skilled labor, $\ell_{mh}$, female skilled labor, $\ell_{fh}$, male unskilled labor, $\ell_{mu}$, and female unskilled labor, $\ell_{fu}$). Building upon the analysis by Fallon and Layard (1975) and Krusell et al. (2000) and the results of Sections 4.2 and 4.3, the sectoral production function is specified as
@@
This production function involves four substitution parameters ($\sigma_{mh}$, $\sigma_{fh}$, $\sigma_{mu}$, and $\sigma_{fu}$) that are less than one and five share parameters ($\theta_{mh}$, $\theta_{fh}$, $\theta_{mu}$, $\theta_{fu}$, and $\theta_{o}$) that lie between zero and one for each sector. The production function exhibits capital–skill complementarity for male labor if $\sigma_{mu}>\sigma_{mh}$ and for female labor if $\sigma_{fu}>\sigma_{fh}$, and capital–gender complementarity for skilled labor if $\sigma_{mh}>\sigma_{fh}$ and for unskilled labor if $\sigma_{mu}>\sigma_{fu}$. We allow for different degrees of capital–skill (capital–gender) complementarity for male and female labor (skilled and unskilled labor) and different degrees of capital–skill–gender complementarity across sectors.
There are three remarks regarding the specifications of the production function. First, ICT capital is located inside the CES function to allow for the possibility that technological change embodied in ICT capital can influence the gender wage gap and the skill wage gap. There were dramatic changes in the price and quantity of ICT equipment (see Figure 4b in the last section for details), whereas there were no such changes in the prices and quantities of non-ICT equipment or structures (see Figure A1 in the appendix for details). Second, ICT capital is placed in the lowest nest in the four-level CES function to allow for the possibility that the four types of labor may be substitutable to varying degrees with ICT capital. Otherwise, the rate of change in the gender (skill) wage gap due to progress in ICT would be the same for skilled and unskilled (male and female) workers, which is not consistent with the data. Finally, the production function is required to be concave (i.e., the substitution parameters must be less than one) to satisfy the economic principle that the relative price of skills should decrease with a rise in the relative supply of skills. For this reason, skilled labor is placed in a nest lower than that of unskilled labor, as in Fallon and Layard (1975) and Krusell et al. (2000), and female labor is placed in a nest lower than that of male labor for each skill type. Appendices A.2.1 and A.2.2 provide further details on the assumptions, implications, and relevance of alternative specifications as well as the robustness of parameter estimates across different specifications. Our specification is consistent with key features of the data described in Sections 4.2 and 4.3, whereas other specifications are not.
@@ 5.2 Identification
Cost minimization implies that the marginal rate of technical substitution equals the ratio of input prices irrespective of the degree of competition in the product market. For the purpose of estimation, we make use of five conditions, concerning (i) the relative price of female skilled labor to ICT capital, (ii) the skilled gender wage gap, (iii) the female skill wage gap, (iv) the unskilled gender wage gap, and (v) the male skill wage gap. The conditions lead to the first-difference equations:
@@
where the error terms, $\Delta v$, are added to account for measurement errors and shocks to non-neutral disembodied technologies, as detailed in Appendix A.2.3, and the log changes of the CES aggregates are approximated by the weighted average of the log changes of factor inputs. It can be shown by Taylor expansion that
@@
where the weights can be calculated directly from the data as $\varphi_{fh,n}=w_{fh,n}\ell_{fh,n}/(r_{i,n}k_{i,n}+w_{fh,n}\ell_{fh,n})$, $\varphi_{mh,n}=w_{mh,n}\ell_{mh,n}/(r_{i,n}k_{i,n}+w_{fh,n}\ell_{fh,n}+w_{mh,n}\ell_{mh,n})$, and $\varphi_{fu,n}=w_{fu,n}\ell_{fu,n}/(r_{i,n}k_{i,n}+w_{fh,n}\ell_{fh,n}+w_{mh,n}\ell_{mh,n}+w_{fu,n}\ell_{fu,n})$. [^6] As a result of the approximations, the share parameters are removed from the system of equations (32)–(36). The substitution parameters alone can be estimated using equations (32)–(36). The share parameters need not be estimated in our analysis because they are not necessary to compute the aggregate elasticities of substitution. Alternatively, the share and substitution parameters can be estimated sequentially using the levels and log differences of the marginal-rate-of-technical substitution equations. We confirm that the approximations above do not affect the estimates of the substitution parameters by comparing the results obtained from the two approaches. Because all time-invariant factors are eliminated by first differencing, the estimates of the system of equations (32)–(36) are immune to the omitted variable bias due to unobserved country-specific characteristics. Differences are taken over five years to avoid the attenuation bias due to measurement error.
Supply shifters are used as instrumental variables to solve the simultaneity bias due to the joint determination of prices and quantities in markets. The instruments used are the lagged population size of each demographic group and the lagged price of ICT investment relative to consumption. This means that we exploit the variation in past birth rates and university enrollment capacity or the variation in investment-specific technology across countries over time. With either of the instruments, the substitution parameters $\sigma_{fh}$, $\sigma_{mh}$, $\sigma_{fu}$, and $\sigma_{mu}$ can be identified sequentially from equations (32)–(36) in the order of those in equations (31), (30), (29), and (28) from the lowest to the highest nest. The error terms are presumably correlated across equations (32)–(36). The substitution parameters in the system of equations (32)–(36) are estimated jointly by the generalized method of moments (GMM) to improve efficiency. Standard errors are clustered at the country level to account for heteroscedasticity and serial correlation. In addition, we control for other factors such as labor market institutions, disembodied factor-biased technological change, and markups in the subsequent analyses.
@@ 5.3 Aggregation
where the sum of the weights on the substitution parameters adds up to one (i.e., $\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{fh}+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{mh}+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{fu}+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{mu}+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{o}+\sum_{n=1}^{N}\pi_{\ell_{f}\ell_{g},n}^{c}=1$). The weights on the substitution parameters depend on the income shares of each factor, as detailed in Appendix A.2.4. The elasticity of substitution in consumption is set at $0.351/(1-\eta)=$ 0.35 in the subsequent analyses, based on the results of Comin, Lashkari and Mestieri (2021). [^7] Comin, Lashkari and Mestieri (2021) estimate the elasticity of substitution among agricultural goods, manufacturing goods, and services using household expenditure data from the United States and aggregate data from a panel of 39 countries. The estimated elasticities of substitution are 0.33 in the United States and 0.35 in OECD countries.
After computing the aggregate elasticities of substitution using the estimated substitution parameters, we can evaluate the effects of technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of the four types of labor. The direction and magnitude of the effects of technological change embodied in ICT capital depend on the presence and degree of capital–skill or capital–gender complementarity (i.e., the sign and magnitude of differences between the aggregate elasticities, $\epsilon_{\ell_{f}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{g}k_{i}}^{\mathcal{C}}$) and the rate of progress in ICT. The degrees of capital–skill and capital–gender complementarities are approximately proportional to the differences in the substitution parameters and the income share of ICT capital relative to the income shares of each type of labor. Appendix A.2.5 provides the exact expressions for the effects of technological change on the relative wages and income shares of the four types of labor.
@@ 6 Results
We start this section by presenting the estimates of sectoral production function parameters and the robustness of the estimates. We then present the estimates of the aggregate elasticities of substitution and the contributions of factor inputs to the relative wages and income shares of the four types of labor. We end this section by discussing the effects of technological change.
@@ 6.1.1 Production function estimates
Table 1 presents the GMM estimates of the substitution parameters in the production functions for manufacturing, utilities, construction and services sectors. Significant differences are observed in the estimates of the four substitution parameters. For manufacturing, utilities, and services sectors, the null hypotheses of $\sigma_{fh}=\sigma_{mh}$, $\sigma_{mh}=\sigma_{fu}$, and $\sigma_{fu}=\sigma_{mu}$ are rejected at the 2 percent significance level (see Panel D of Table A1 in the appendix for details). The estimates suggest that ICT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor (i.e., $\sigma_{fh}<\sigma_{mh}<\sigma_{fu}<\sigma_{mu}$). Moreover, the null hypothesis that the substitution parameters are equal across the three sectors is rejected at the 3 percent significance level. For the construction sector, the substitution parameters are estimated under the null hypothesis of equal parameters (i.e., $\sigma_{fh}=\sigma_{mh}=\sigma_{fu}=\sigma_{mu}$) because the null hypothesis cannot be rejected at the 39 percent significance level. The estimates of the substitution parameters differ significantly across sectors. The estimates remain similar regardless of whether the lagged population size or the lagged investment price is used as an instrument. The estimates with the lagged population size are used in the subsequent analyses because they are more precise.
@@
Manufacturing –0.698 0.287 0.602 0.782 –0.780 0.136 0.540 0.693
(0.337) (0.047) (0.043) (0.056) (0.303) (0.074) (0.069) (0.052)
Utilities –0.904 0.227 0.427 0.657 –1.076 0.064 0.329 0.551
(0.289) (0.083) (0.057) (0.059) (0.324) (0.158) (0.084) (0.051)
Construction 0.633 0.633 0.633 0.633 0.506 0.506 0.506 0.506
(0.083) (0.083) (0.083) (0.083) (0.092) (0.092) (0.092) (0.092)
Services –0.415 0.342 0.599 0.871 –0.456 0.203 0.561 0.789
(0.136) (0.070) (0.034) (0.049) (0.119) (0.075) (0.036) (0.058)
Notes: Standard errors in parentheses are clustered at the country level. Lagged population size and lagged ICT investment price are used as instrumental variables in the first to fourth columns and the fifth to the last columns, respectively. The production function parameters for the construction sector are estimated under the restriction of $\sigma_{fh}=\sigma_{mh}=\sigma_{fu}=\sigma_{mu}$ because the restriction cannot be rejected.

### v9

@@ 4.3 Factor substitution
Manufacturing 0.765 1.048 0.880 0.965 0.576 0.986 1.298 1.296 0.548 0.891 1.180 1.183
(0.152) (0.164) (0.185) (0.160) {0.415, 0.763} {0.872, 1.144} {1.104, 1.579} {1.134, 1.531} {0.401, 0.700} {0.753, 1.050} {0.955, 1.461} {0.990, 1.415}
Utilities 0.281 0.735 0.656 0.968 0.440 0.950 1.118 1.260 0.457 0.877 1.037 1.157
(0.278) (0.113) (0.160) (0.161) {0.293, 0.630} {0.700, 1.287} {0.895, 1.435} {1.034, 1.587} {0.319, 0.629} {0.603, 1.242} {0.775, 1.386} {0.891, 1.503}
Construction 0.637 0.832 0.775 0.663 1.066 1.414 1.458 1.514 0.888 1.176 1.289 1.342
(0.328) (0.447) (0.143) (0.169) {0.440, 2.093} {0.798, 2.335} {0.885, 2.325} {0.973, 2.345} {0.488, 1.508} {0.749, 1.789} {0.848, 1.932} {0.920, 1.971}
Services 0.691 0.905 0.788 0.920 0.727 0.977 1.208 1.298 0.689 0.890 1.104 1.167
(0.189) (0.100) (0.094) (0.079) {0.580, 0.885} {0.849, 1.153} {1.057, 1.418} {1.137, 1.543} {0.564, 0.814} {0.772, 1.019} {0.978, 1.249} {1.035, 1.322}
Notes: Estimates of $\epsilon_{f,n}$ in equation (23) for female skilled labor ($f=fh$), male skilled labor ($f=mh$), female unskilled labor ($f=fu$), and male unskilled labor ($f=mu$) are reported. Standard errors are in parentheses of the first four columns. The 95% Anderson–Rubin confidence sets are in curly brackets of the last eight columns. Both of them are robust to clustering at the country level. Lagged population size and lagged ICCT investment price are used as instrumental variables in the fifth to eighth columns and the ninth to the last columns, respectively.
Table 1 presents the ordinary least squares (OLS) and first-difference generalized method of moments (GMM) estimates of $\varepsilon_{f,n}$ for female skilled labor ($f=fh$), male skilled labor ($f=mh$), female unskilled labor ($f=fu$), and male unskilled labor ($f=mu$) in manufacturing, utilities, construction, and services sectors. Both estimates indicate that the degrees of substitution between ICCT capital and the four types of labor are smaller in the order of female skilled, male skilled, female unskilled, and male unskilled labor for all sectors (i.e., $\varepsilon_{fh,n}<\varepsilon_{mh,n}<\varepsilon_{fu,n}<\varepsilon_{mu,n}$ for all $n$). The first-difference GMM estimates differ by the type of labor more than the OLS estimates. In both cases, the differences in the estimates by the type of labor are more significant for manufacturing, utilities, and services sectors than for the construction sector. Similar results are obtained for the 1980–1995 and 1990–2005 periods. Although the degrees of substitution between ICCT capital and unskilled labor become smaller from the former to the latter period, ICCT capital is more substitutable with male and female unskilled labor than with male and female skilled labor. The results described thus far are reinforced by the Anderson and Rubin (1949) confidence sets robust to weak instruments. The first-difference GMM estimates remain largely unchanged regardless of whether the lagged population size or the lagged investment price is used as an instrument. Overall, ICCT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor.
@@ 5 Estimation
This section describes the specification, identification, estimation, and aggregation of the sectoral production function. We consider the sectoral production function in which all four types of labor are imperfect substitutes for one another and substitutable to varying degrees with ICCT capital.
@@ 5.1 Specification
The economy consists of multiple sectors. In each sector, production requires two types of capital (ICCT capital, $k_{i}$, and non-ICCT capital, $k_{o}$) and four types of labor (male skilled labor, $\ell_{mh}$, female skilled labor, $\ell_{fh}$, male unskilled labor, $\ell_{mu}$, and female unskilled labor, $\ell_{fu}$). Building upon the analysis by Fallon and Layard (1975) and Krusell et al. (2000) and in Sections 4.2 and 4.3, the sectoral production function is specified as
@@
This production function involves four substitution parameters ($\sigma_{mh}$, $\sigma_{fh}$, $\sigma_{mu}$, and $\sigma_{fu}$) that are less than one and five share parameters ($\theta_{mh}$, $\theta_{fh}$, $\theta_{mu}$, $\theta_{fu}$, and $\theta_{o}$) that lie between zero and one for each sector. The production function exhibits capital–skill complementarity for male labor if $\sigma_{mu}>\sigma_{mh}$ and for female labor if $\sigma_{fu}>\sigma_{fh}$, and capital–gender complementarity for skilled labor if $\sigma_{mh}>\sigma_{fh}$ and for unskilled labor if $\sigma_{mu}>\sigma_{fu}$. We allow for different degrees of capital–skill complementarity for male and female labor and different degrees of capital–gender complementarity for skilled and unskilled labor. In addition, we allow for different degrees of capital–skill–gender complementarity across sectors.
The multi-level nested CES production function (24) is not only analytically tractable enough to derive the general equilibrium effects of technological change on the relative wages and income shares of different types of labor but also flexible enough to allow for different degrees of capital–skill and capital–gender complementarities. There are three features in the specification of the production function. First, ICCT capital is located inside the CES function to allow for the possibility that technological change embodied in ICCT capital can influence the gender wage gap and the skill wage gap. There were dramatic changes in the price and quantity of ICCT equipment (see Figure 4b in the last section), whereas there were no such changes in the prices and quantities of non-ICCT equipment or structures (see Figure A1 in Appendix D.3). Second, ICCT capital is placed in the lowest nest in the four-level CES function to allow for the possibility that the four types of labor may be substitutable to varying degrees with ICCT capital. Otherwise, the rate of change in the gender (skill) wage gap due to progress in ICCT would be the same for skilled and unskilled (male and female) workers, which is not consistent with the data. Finally, the production function is required to be concave (i.e., the substitution parameters must be less than one) to satisfy the economic principle that the relative price of skills should decrease with a rise in the relative supply of skills. For this reason, skilled labor is placed in a nest lower than that of unskilled labor, as in Fallon and Layard (1975) and Krusell et al. (2000), and female labor is placed in a nest lower than that of male labor for each skill type. Our specification is consistent with key features of the data described in Sections 4.2 and 4.3. Appendices E.1 and E.2 provide further details on the assumptions, implications, and relevance of alternative specifications.
@@ 5.2 Identification
Cost minimization implies that the marginal rate of technical substitution equals the ratio of input prices irrespective of the degree of competition in goods markets. For the purpose of our analysis, we make use of five conditions, concerning (i) the relative price of female skilled labor to ICCT capital, (ii) the skilled gender wage gap, (iii) the female skill wage gap, (iv) the unskilled gender wage gap, and (v) the male skill wage gap, to estimate the production function parameters. The conditions lead to the following first-difference equations:
@@
where the error terms, $\Delta v$, are added to account for measurement errors and shocks to non-neutral disembodied technologies, as detailed in Appendix E.3, and the log changes of the CES aggregates are approximated by the weighted average of the log changes of factor inputs. It can be shown by Taylor expansion that
@@
where the weights can be calculated directly from the data as $\varphi_{fh,n}=w_{fh,n}\ell_{fh,n}/(r_{i,n}k_{i,n}+w_{fh,n}\ell_{fh,n})$, $\varphi_{mh,n}=w_{mh,n}\ell_{mh,n}/(r_{i,n}k_{i,n}+w_{fh,n}\ell_{fh,n}+w_{mh,n}\ell_{mh,n})$, and $\varphi_{fu,n}=w_{fu,n}\ell_{fu,n}/(r_{i,n}k_{i,n}+w_{fh,n}\ell_{fh,n}+w_{mh,n}\ell_{mh,n}+w_{fu,n}\ell_{fu,n})$. [^12] As a result of these approximations, the share parameters are eliminated from the system of equations (28)–(32). The substitution parameters can be estimated solely using equations (28)–(32). However, it is unnecessary to estimate the share parameters in our analysis because they are not required to compute the aggregate elasticities of substitution. Alternatively, the share and substitution parameters can be estimated sequentially using the levels and log differences of the marginal-rate-of-technical substitution equations. We confirm that the approximations do not affect the estimates of the substitution parameters by comparing the results obtained from both approaches. Because all time-invariant factors are eliminated by first differencing, the estimates of the system of equations (28)–(32) are immune to the omitted variable bias due to unobserved country-specific characteristics. Differences are taken over five years to avoid the attenuation bias due to measurement error.
Because prices and quantities are jointly determined in markets, identifying equations (28)–(32) for factor demand requires using supply shifters as instrumental variables. The instruments used are the lagged population size of each demographic group and the lagged price of ICCT investment relative to consumption. This means that we exploit the variation in past birth rates and university enrollment capacity or the variation in investment-specific technology across countries over time by taking advantage of cross-country panel data. With either of the instruments, it is straightforward to see that the substitution parameters $\sigma_{fh}$, $\sigma_{mh}$, $\sigma_{fu}$, and $\sigma_{mu}$ can be identified sequentially from equations (28)–(32) in the order of those in equations (27), (26), (25), and (24) from the lowest to the highest nest. The error terms are presumably correlated across equations (28)–(32). The substitution parameters in the system of equations (28)–(32) are estimated jointly by GMM to improve efficiency. Standard errors are clustered at the country level to account for heteroscedasticity and serial correlation. In addition, we control for other factors such as labor market institutions, disembodied factor-biased technological change, and markups.
@@ 5.3 Aggregation
where the sum of the weights on the substitution parameters adds up to one. The weights on the substitution parameters depend on the income shares of each factor, as detailed in Appendix E.4. The elasticity of substitution in consumption is set at $1/(1-\eta)=0.9$, as in Baqaee and Farhi (2019).
After computing the aggregate elasticities of substitution using the estimated substitution parameters, we can evaluate the effects of technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of the four types of labor. The direction and magnitude of the effects of technological change embodied in ICCT capital depend on the presence and degree of capital–skill or capital–gender complementarity (i.e., the sign and magnitude of differences between the aggregate elasticities, $\epsilon_{\ell_{f}k_{i}}^{\mathcal{C}}-\epsilon_{\ell_{g}k_{i}}^{\mathcal{C}}$) and the rate of progress in ICCT. The degrees of capital–skill and capital–gender complementarities are approximately proportional to the differences in the substitution parameters and the income share of ICCT capital relative to the income shares of each type of labor. Appendix E.5 provides the exact expressions for the effects of technological change on the relative wages and income shares of the four types of labor.
@@ 6 Results
This section begins by presenting the estimates of sectoral production function parameters and assessing the robustness of the estimates. It then reports the estimates of the aggregate elasticities of substitution and the quantitative contributions of factor inputs to the relative wages and income shares of the four types of labor. The section ends with findings on the effects of technological change.
@@ 6.1.1 Production function estimates
Table 2 presents the GMM estimates of the substitution parameters in the production functions for manufacturing, utilities, construction and services sectors. For comparison it also presents corresponding estimates under the assumption that the aggregate production function takes the form of equation (24). Significant differences are observed in the estimates of the four substitution parameters for three out of four sectors. For manufacturing, utilities, and services sectors, the null hypotheses of $\sigma_{fh}=\sigma_{mh}$, $\sigma_{mh}=\sigma_{fu}$, and $\sigma_{fu}=\sigma_{mu}$ are rejected at the 2 percent significance level (see Panel D of Table A1 in Appendix E.1). The estimates suggest that ICCT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor (i.e., $\sigma_{fh}<\sigma_{mh}<\sigma_{fu}<\sigma_{mu}$). Moreover, the null hypothesis that the substitution parameters are equal across the three sectors is rejected at the 2 percent significance level. [^13] The null hypotheses that the substitution parameters are equal between the manufacturing and utilities sectors, between the utilities and services sectors, and between the manufacturing and services sectors are rejected at the 10 percent, 3 percent, and 13 percent significance levels, respectively. For the construction sector, the substitution parameters are estimated under the null hypothesis of equal parameters (i.e., $\sigma_{fh}=\sigma_{mh}=\sigma_{fu}=\sigma_{mu}$) because the null hypothesis cannot be rejected at the 39 percent significance level. [^14] On average in our sample, the shares of college graduates and women each account for less than 10 percent in the construction sector. This pattern is consistent with the absence of capital–skill and capital–gender complementarities. These results are not an artifact of an arbitrary nesting structure of the CES production function. The relative magnitude of the estimated degrees of factor substitution is consistent with that reported in Table 1. The estimates remain similar regardless of whether the lagged population size or the lagged investment price is used as an instrument. The estimates obtained with the lagged population size are used in the subsequent analyses because they are more precise. When the substitution parameters are estimated under the assumption that the aggregate production function for the non-primary sector exists, the parameter estimates differ from those obtained for the manufacturing, utilities, and construction sectors.
@@
Manufacturing –0. 689 0. 288 0. 600 0. 776 –0. 780 0. 136 0. 540 0. 693
(0. 319) (0. 046) (0. 043) (0. 054) (0. 303) (0. 074) (0. 069) (0. 052)
Utilities –0. 962 0. 184 0. 404 0. 631 –1. 076 0. 064 0. 329 0. 551
(0. 362) (0. 104) (0. 063) (0. 058) (0. 324) (0. 158) (0. 084) (0. 051)
Construction 0. 628 0. 628 0. 628 0. 628 0. 506 0. 506 0. 506 0. 506
(0. 077) (0. 077) (0. 077) (0. 077) (0. 092) (0. 092) (0. 092) (0. 092)
Services –0. 413 0. 340 0. 598 0. 861 –0. 456 0. 203 0. 561 0. 789
(0. 134) (0. 067) (0. 034) (0. 048) (0. 119) (0. 075) (0. 036) (0. 058)
Non-primary –0. 477 0. 333 0. 604 0. 847 –0. 517 0. 193 0. 560 0. 784
(0. 171) (0. 061) (0. 036) (0. 047) (0. 145) (0. 071) (0. 041) (0. 044)
Notes: Standard errors in parentheses are clustered at the country level. Lagged population size and lagged ICCT investment price are used as instrumental variables in the first to fourth columns and the fifth to the last columns, respectively. The production function parameters for the construction sector are estimated under the restriction of $\sigma_{fh}=\sigma_{mh}=\sigma_{fu}=\sigma_{mu}$ because the restriction cannot be rejected.